OpenStudy (anonymous):

Given the point (3, 4) and the slope of 6, find y when x = 27 help!!

zepdrix (zepdrix):

Sorry my OpenStudy crashed :( I need to retype all of that. Grr

zepdrix (zepdrix):

We can use the Point-Slope form of a line:\[\Large y-y_o=m(x-x_o)\] We were given the slope:\[\Large m=6\]And a starting point,\[\Large (x_o,y_o)=(3,4)\]And an x value,\[\Large x=27\]

zepdrix (zepdrix):

We can plug all those pieces in and solve for y! :O

OpenStudy (anonymous):

so y= 6x+4?

zepdrix (zepdrix):

Hmm you didn't plug in the x=27. And it looks like you did something wrong in there :[ Let's see... Plugging in our values gives us,\[\Large y-y_o=m(x-x_o) \quad\color{orangered}{\to}\quad y-4=6(27-3)\]

OpenStudy (anonymous):

y= 163?

zepdrix (zepdrix):

27 minus 3, \[\large y-4=6(24)\] 6 times 24,\[\large y-4=144\]Add 4 to each side,\[\large y=144+4\]